OpenStudy (mortonsalt):

Complex numbers! (1+i)^14?

satellite73 (satellite73):

job one is to write it in trig form (polar form) do you know how to do that?

satellite73 (satellite73):

i.e. write it as \[r\left(\cos(\theta)+i\sin(\theta)\right)\]

OpenStudy (mortonsalt):

sqrt 2 (cos pi/4 + i sin pi/4)?

satellite73 (satellite73):

yes!

OpenStudy (mortonsalt):

@Satellite73 Thank you! Just a massive brain fart.

satellite73 (satellite73):

so you have only two easy steps now \[\sqrt{2}^{14}\] for the absolute value (modulus) and \[14\times \frac{\pi}{4}\]for the angle

satellite73 (satellite73):

is it clear what \[\sqrt{2}^{14}\] is ?

satellite73 (satellite73):

without a calculator i mean with a calculator it is easy

OpenStudy (mortonsalt):

Yep! Just that one step that really bugged me. Thanks! @Satellite73

satellite73 (satellite73):

yw